ABC Corp. has sales of $24m in 2017. Its ROE was 12% with a total asset turnover of 3.0X. Its Debt/Equity ratio is 2/3. Find its Net Income:

Select one:

a. $720,000

b. None of the Above

c. $560,000

d. $576,000

e. $420,000
Answer #1

Debt / Equity = 2/3, Therefore, Debt / Assets = 2/5 and Equity / Assets = 3/5

Total Asset Turnover = Sales / Assets = 3 and ROE = Net Profit / Shareholder's Equity = 0.12 or 12 %

Total Asset Turnover / ROE = (Sales / Assets ) x (Equity / Net Income)

3 / 0.12 = (Sales / Net Income) x (Equity / Assets)

(3 / 0.12) x (5 / 3 ) = (24 / Net Income)

Net Income = 24 x (3/5) x (0.12 / 3) = 24 x 0.04 x 0.6 = $ 0.576 million or $ 5760000

Hence, the correct option is (d)